# Ego-centric Predictive Model Conditioned on Hand Trajectories

### v1

@@ Abstract
In egocentric scenarios, anticipating both the next action and its visual outcome is essential for understanding human–object interactions and for enabling robotic planning. However, existing paradigms fall short of jointly modeling these aspects. Vision-Language-Action (VLA) models focus on action prediction but lack explicit modeling of how actions influence the visual scene, while video prediction models generate future frames without conditioning on specific actions—often resulting in implausible or contextually inconsistent outcomes. To bridge this gap, we propose a unified two-stage predictive framework that jointly models action and visual future in egocentric scenarios, conditioned on hand trajectories. In the first stage, we perform consecutive state modeling to process heterogeneous inputs—visual observations, language, and action history—and explicitly predict future hand trajectories. In the second stage, we introduce causal cross-attention to fuse multi-modal cues, leveraging inferred action signals to guide an image-based Latent Diffusion Model (LDM) for frame-by-frame future video generation. Our approach is the first unified model designed to handle both egocentric human activity understanding and robotic manipulation tasks, providing explicit predictions of both upcoming actions and their visual consequences. Extensive experiments on Ego4D, BridgeData, and RLBench demonstrate that our method outperforms state-of-the-art baselines in both action prediction and future video synthesis. Code is available at https://github.com/binjiezhang/Ego-PM/tree/main.
@@ Conclusion
Our contributions include introducing a simple yet powerful conditioning signal (hand position annotations) to guide the learning of action–effect relationships, and a consecutive state modeling strategy that effectively integrates heterogeneous inputs. Crucially, our model operates without requiring any manual action annotations at test time, making it practical for real-world deployment. Extensive experiments on three distinct datasets – Ego4D (human egocentric videos), BridgeData (human-to-robot scenarios), and RLBench (robotic manipulation tasks) – validate the effectiveness of our model. It consistently outperforms state-of-the-art baselines in both the accuracy of predicted actions and the quality of generated future video sequences. These results underscore the generality of our approach, as the same model can be seamlessly applied to both human and robotic perspectives.
Future work can build upon this foundation by exploring even richer conditioning signals and by extending the model to handle longer-horizon predictions and more complex multi-agent interactions. Overall, our hand-conditioned egocentric predictive model opens new avenues for intuitive and interactive AI systems that understand and predict the coupling between actions and their effects in the world.

### v2

@@ Abstract
In egocentric scenarios, anticipating both the next action and its visual outcome is essential for understanding human–object interactions and for enabling robotic planning. However, existing paradigms fall short of jointly modeling these aspects. Vision-Language-Action (VLA) models focus on action prediction but lack explicit modeling of how actions influence the visual scene, while video prediction models generate future frames without conditioning on specific actions—often resulting in implausible or contextually inconsistent outcomes. To bridge this gap, we propose a unified two-stage predictive framework that jointly models action and visual future in egocentric scenarios, conditioned on hand trajectories. In the first stage, we perform consecutive state modeling to process heterogeneous inputs—visual observations, language, and action history—and explicitly predict future hand trajectories. In the second stage, we introduce causal cross-attention to fuse multi-modal cues, leveraging inferred action signals to guide an image-based Latent Diffusion Model (LDM) for frame-by-frame future video generation. Our approach is the first unified model designed to handle both egocentric human activity understanding and robotic manipulation tasks, providing explicit predictions of both upcoming actions and their visual consequences. Extensive experiments on Ego4D, BridgeData, and RLBench demonstrate that our method outperforms state-of-the-art baselines in both action prediction and future video synthesis. Code is available at https://github.com/showlab/Ego-PM.
@@ Conclusion
Our contributions include introducing a simple yet powerful conditioning signal (hand position annotations) to guide the learning of action–effect relationships, and a consecutive state modeling strategy that effectively integrates heterogeneous inputs. Crucially, our model operates without requiring any manual action annotations at test time, making it practical for real-world deployment. Extensive experiments on three distinct datasets – Ego4D (human egocentric videos), BridgeData (human-to-robot scenarios), and RLBench (robotic manipulation tasks) – validate the effectiveness of our model. It consistently outperforms state-of-the-art baselines in both the accuracy of predicted actions and the quality of generated future video sequences. These results underscore the generality of our approach, as the same model can be seamlessly applied to both human and robotic perspectives. 6—— Future work can build upon this foundation by exploring even richer conditioning signals and by extending the model to handle longer-horizon predictions and more complex multi-agent interactions. Overall, our hand-conditioned egocentric predictive model opens new avenues for intuitive and interactive AI systems that understand and predict the coupling between actions and their effects in the world.
Appendix I: Related Work
In the main paper, we briefly summarize relevant research on text-based video generation, robotics and vision-language-action (VLA) models, and egocentric vision-language models. In addition to these areas, we also provide a discussion of related world model approaches.
World Models.
World models (Ha and Schmidhuber 2018) enable intelligent agents to understand, predict, and make decisions based on interactions with their environments. In computer vision, world models are widely used for video prediction (Lotter, Kreiman, and Cox 2017; Lee, Strobel, and Finn 2018), which is critical for downstream tasks such as autonomous driving (Hu et al. 2023; Shi et al. 2023; Zheng et al. 2025; Tian et al. 2024) and robotics (Liang et al. 2022; Lu et al. 2025; Firoozi et al. 2023). These models generate plausible future frames by conditioning on the agent’s current state, actions, and environmental context, thereby enhancing temporal coherence and action prediction accuracy.
LWM (Liu et al. 2024) tackles long-sequence prediction via masked sequence packing to improve training efficiency. However, it directly concatenates video frames without explicitly modeling consecutive states. GAIA-1 (Hu et al. 2023) predicts future frames conditioned on visual inputs and action sequences, facilitating real-time event anticipation in autonomous driving. Despite their effectiveness, GAIA-1 and related models (Wang et al. 2024b) primarily operate from exocentric (third-person) viewpoints, which limits their capacity to model fine-grained egocentric details such as hand movements. While world models have made considerable progress, they still face challenges in egocentric and interactive tasks where a user-centered, first-person view is essential for immersive applications.
Appendix II: Technical Details
Overview of Ego-PM.
An overview of the Ego-Centric Predictive Model (Ego-PM) is illustrated in Fig. 4, inspired by the architecture of the original World Model (Ha and Schmidhuber 2018). The model consists of three core components:
  • Multimodal Encoders extract features from visual, textual, and action inputs. Incorporating the action space is critical for accurate prediction.
  • An Autoregressive Model learns cross-modal correlations to enhance temporal coherence and explicitly predict future hand trajectories.
  • A Controller uses these representations to predict future frames, generate narrative descriptions, and anticipate upcoming actions.
Figure 4: Overview of Ego-PM architecture, comprising multimodal encoders, an autoregressive model, and controllers. The model predicts consecutive future states, including both frames and actions. Darker colors denote input states, and lighter colors indicate predicted states.
Training Details.
We use Bi-State Prediction to illustrate the CoSMo (Consecutive State Modeling) strategy. Given triplets of input data $\{\mathcal{V}_{t:1\rightarrow 2},\mathcal{A}_{t:1\rightarrow 2},\mathcal{T}\}$, the model learns to predict future states from consecutive inputs. We introduce a [MASK] token to represent a zero-initialized state at $t=0$. The model first predicts the state at $t=2$ using inputs from $t=0$ and $t=1$, then continues with $t=1$ and $t=2$ to predict $t=3$, and so on.
Model Input Modality Prediction Capability
Scene Text Action Frame Text Action
LWM Exocentric
GAIA-1 Exocentric Vehicle behavior
WorldDreamer Exocentric Vehicle behavior
LEGO Egocentric
OpenVLA Robot
PEVA Human Body skeleton
Ours (Ego-PM) Omni Hand pose / Robot kinematics
Table 6: Comparison of baselines with respect to input modality and prediction capabilities. Our model uniquely supports egocentric inputs with multimodal conditions (text, visual, action) and predicts future frames, narrative text, and action movements.
Model Frame Prediction
[…]
PVEA-B (original, DiT-based) 74.338 16.01 33.7
PVEA-B (reproduced) 78.56 15.61 35.8
Ours (LDM-based) 76.21 16.10 33.4
Table 7: Frame prediction results on the Nymeria dataset. Lower FID and LPIPS, and higher PSNR indicate better performance.
Comparison of Model Capabilities.
Table 6 compares various models in terms of input modality and prediction ability. LWM (Liu et al. 2024) can process long exocentric videos and generate both frames and text. GAIA-1 (Hu et al. 2023) and WorldDreamer (Wang et al. 2024b) use exocentric views and vehicle behavior for frame prediction. LEGO (Lai et al. 2025), though not a full world model due to missing action prediction, remains a strong baseline for egocentric frame generation.
Pseudo-code for Consecutive State Modeling (CoSMo)
To improve temporal consistency and action prediction in egocentric settings, we introduce CoSMo, which utilizes both current and previous states to predict future ones. Unlike prior work (Lai et al. 2025) that relies on only the current state, our method captures richer temporal dependencies.
Appendix III: Additional Experiments
Human-Skeleton-Based Video Prediction.
Beyond hand trajectory modeling, we explore the use of full-body motion as an action representation. Specifically, we replace the original $4\times 2$ hand coordinates with 17 human keypoints (skeleton) to examine whether explicitly modeling richer action space improves video generation.
We use the Nymeria dataset (Ma et al. 2024), which contains egocentric RGB frames, 3D body pose annotations, and textual descriptions in real-world scenarios. We use PEVA (Bai et al. 2025) as the baseline and filter training samples with available text.
PEVA is a video-based DiT method that requires 3–15 continuous frames and corresponding skeleton inputs to predict 64 future frames. In contrast, our method (LDM-based) only requires two frames and a text prompt. We first predict the action representation, then generate future video frames iteratively using action embeddings.
Results in Table 7 show that our method outperforms the PEVA baseline using the same training set, demonstrating that explicit action modeling enhances prediction quality.
Algorithm 1 Consecutive State Modeling (CoSMo)
1:Input: Video frames $\mathcal{V}_{t:1\rightarrow 2}$, Actions $\mathcal{A}_{t:1\rightarrow 2}$, Text prompt $\mathcal{T}$
2:Output: Predicted frames $\hat{\mathcal{V}}_{t:>2}$, Actions $\hat{\mathcal{A}}_{t:>2}$, Action description $\mathcal{T}_{\mathcal{A}}$
3:Initialize masked inputs: $v_{0}\leftarrow[\text{MASK}],a_{0}\leftarrow[\text{MASK}]$
4:for each training step in Stage I do
5:  Sample $\{\mathcal{V}_{t:1\rightarrow 2},\mathcal{A}_{t:1\rightarrow 2},\mathcal{T}\}$
6:  Predict $\{\hat{v}_{2},\hat{a}_{2},\mathcal{T}_{\mathcal{A}}\}$ using $(v_{0},v_{1},a_{0},a_{1},\mathcal{T})$
7:  Predict $\{\hat{v}_{3},\hat{a}_{3}\}$ using $(v_{1},\hat{v}_{2},a_{1},\hat{a}_{2},\mathcal{T})$
8:  Compute loss $\mathcal{L}_{\text{CoSMo}}$ and update model parameters
9:end for
10:for each training step in Stage II do
11:  Sample $\{\mathcal{V}_{t:1\rightarrow 2},\mathcal{A}_{t:1\rightarrow 2},\mathcal{T},\mathcal{T}_{\mathcal{A}}\}$
12:  Predict $\hat{v}_{2}$ using $(v_{0},v_{1},a_{0},a_{1},\mathcal{T},\mathcal{T}_{\mathcal{A}})$
13:  Predict $\hat{v}_{3}$ using $(v_{1},\hat{v}_{2},a_{1},\hat{a}_{2},\mathcal{T},\mathcal{T}_{\mathcal{A}})$
14:  Compute loss $\mathcal{L}_{\text{LDM}}$ and update model parameters
15:end for